$SPAD/src/input derham.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool derham.output
)set message test on
)set message auto off
)clear all
--S 1 of 33
coefRing := Integer
--R
--R
--R   (1)  Integer
--R                                                                 Type: Domain
--E 1
--S 2 of 33
lv : List Symbol := [x,y,z]
--R
--R
--R   (2)  [x,y,z]
--R                                                           Type: List(Symbol)
--E 2
--S 3 of 33
der := DERHAM(coefRing,lv)
--R
--R
--R   (3)  DeRhamComplex(Integer,[x,y,z])
--R                                                                 Type: Domain
--E 3
--S 4 of 33
R := Expression coefRing
--R
--R
--R   (4)  Expression(Integer)
--R                                                                 Type: Domain
--E 4
--S 5 of 33
f : R := x**2*y*z-5*x**3*y**2*z**5

--R
--R
--R            3 2 5    2
--R   (5)  - 5x y z  + x y z
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 33
g : R := z**2*y*cos(z)-7*sin(x**3*y**2)*z**2
--R
--R
--R            2     3 2       2
--R   (6)  - 7z sin(x y ) + y z cos(z)
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 33
h : R :=x*y*z-2*x**3*y*z**2
--R
--R
--R            3   2
--R   (7)  - 2x y z  + x y z
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 33
dx : der := generator(1)
--R
--R
--R   (8)  dx
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 8
--S 9 of 33
dy : der := generator(2)
--R
--R
--R   (9)  dy
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 9
--S 10 of 33
dz : der := generator(3)
--R
--R
--R   (10)  dz
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 10
--S 11 of 33

[dx,dy,dz] := [generator(i)$der for i in 1..3]
--R
--R
--R   (11)  [dx,dy,dz]
--R                                   Type: List(DeRhamComplex(Integer,[x,y,z]))
--E 11
--S 12 of 33
alpha : der := f*dx + g*dy + h*dz
--R
--R
--R   (12)
--R          3   2                   2     3 2       2
--R     (- 2x y z  + x y z)dz + (- 7z sin(x y ) + y z cos(z))dy
--R   +
--R          3 2 5    2
--R     (- 5x y z  + x y z)dx
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 12
--S 13 of 33
beta  : der := cos(tan(x*y*z)+x*y*z)*dx + x*dy
--R
--R
--R   (13)  x dy + cos(tan(x y z) + x y z)dx
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 13
--S 14 of 33
exteriorDifferential alpha;
--R
--R
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 14
--S 15 of 33
exteriorDifferential %
--R
--R
--R   (15)  0
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 15
--S 16 of 33
gamma := alpha * beta
--R
--R
--R   (16)
--R        4   2    2               3   2
--R     (2x y z  - x y z)dy dz + (2x y z  - x y z)cos(tan(x y z) + x y z)dx dz

--R   +
--R       2     3 2       2                                   4 2 5    3
--R   ((7z sin(x y ) - y z cos(z))cos(tan(x y z) + x y z) - 5x y z  + x y z)dx dy
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 16
--S 17 of 33
exteriorDifferential(gamma) - (exteriorDifferential(alpha)*beta - _
                                alpha * exteriorDifferential(beta))
--R
--R
--R   (17)  0
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 17
--S 18 of 33
a : BOP := operator('a)
--R
--R
--R   (18)  a
--R                                                          Type: BasicOperator
--E 18
--S 19 of 33
b : BOP := operator('b)
--R
--R
--R   (19)  b
--R                                                          Type: BasicOperator
--E 19
--S 20 of 33
c : BOP := operator('c)
--R
--R
--R   (20)  c
--R                                                          Type: BasicOperator
--E 20
--S 21 of 33
sigma := a(x,y,z) * dx + b(x,y,z) * dy + c(x,y,z) * dz
--R
--R
--R   (21)  c(x,y,z)dz + b(x,y,z)dy + a(x,y,z)dx
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 21
--S 22 of 33
theta  := a(x,y,z) * dx * dy + b(x,y,z) * dx * dz + c(x,y,z) * dy * dz
--R

--R
--R   (22)  c(x,y,z)dy dz + b(x,y,z)dx dz + a(x,y,z)dx dy
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 22
--S 23 of 33
totalDifferential(a(x,y,z))$der
--R
--R
--R   (23)  a  (x,y,z)dz + a  (x,y,z)dy + a  (x,y,z)dx
--R          ,3             ,2             ,1
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 23
--S 24 of 33
exteriorDifferential sigma
--R
--R
--R   (24)
--R     (c  (x,y,z) - b  (x,y,z))dy dz + (c  (x,y,z) - a  (x,y,z))dx dz
--R       ,2           ,3                  ,1           ,3
--R   +
--R     (b  (x,y,z) - a  (x,y,z))dx dy
--R       ,1           ,2
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 24
--S 25 of 33
exteriorDifferential theta
--R
--R
--R   (25)  (c  (x,y,z) - b  (x,y,z) + a  (x,y,z))dx dy dz
--R           ,1           ,2           ,3
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 25
--S 26 of 33
one : der := 1
--R
--R
--R   (26)  1
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 26
--S 27 of 33
g1 : der := a([x,t,y,u,v,z,e]) * one
--R
--R
--R   (27)  a(x,t,y,u,v,z,e)
--R                                         Type: DeRhamComplex(Integer,[x,y,z])

--E 27
--S 28 of 33
h1 : der := a([x,y,x,t,x,z,y,r,u,x]) * one
--R
--R
--R   (28)  a(x,y,x,t,x,z,y,r,u,x)
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 28
--S 29 of 33
exteriorDifferential g1
--R
--R
--R   (29)  a  (x,t,y,u,v,z,e)dz + a  (x,t,y,u,v,z,e)dy + a  (x,t,y,u,v,z,e)dx
--R          ,6                     ,3                     ,1
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 29
--S 30 of 33
exteriorDifferential h1
--R
--R
--R   (30)
--R     a  (x,y,x,t,x,z,y,r,u,x)dz
--R      ,6
--R   +
--R     (a  (x,y,x,t,x,z,y,r,u,x) + a  (x,y,x,t,x,z,y,r,u,x))dy
--R       ,7                         ,2
--R   +
--R         a   (x,y,x,t,x,z,y,r,u,x) + a  (x,y,x,t,x,z,y,r,u,x)
--R          ,10                         ,5
--R       +
--R         a  (x,y,x,t,x,z,y,r,u,x) + a  (x,y,x,t,x,z,y,r,u,x)
--R          ,3                         ,1
--R    *
--R       dx
--R                                         Type: DeRhamComplex(Integer,[x,y,z])
--E 30
--S 31 of 33
coefficient(gamma, dx*dy)
--R
--R
--R            2     3 2       2                                   4 2 5    3
--R   (31)  (7z sin(x y ) - y z cos(z))cos(tan(x y z) + x y z) - 5x y z  + x y z
--R                                                    Type: Expression(Integer)
--E 31
--S 32 of 33

coefficient(gamma, one)
--R
--R
--R   (32)  0
--R                                                    Type: Expression(Integer)
--E 32
--S 33 of 33
coefficient(g1,one)
--R
--R
--R   (33)  a(x,t,y,u,v,z,e)
--R                                                    Type: Expression(Integer)
--E 33
)spool
)lisp (bye)

———-

References^{[1] nothing}

10